OpenStudy (anonymous):

find f^-1(x) f(x)=x^2-4, x>_0

OpenStudy (alexwee123):

f^-1(x) is the inverse change around the y with x so x=y^2-3 x-2=y^2 y=(x-2)^1/2

OpenStudy (anonymous):

lol i like how the 4 changed to a 3 and then to a -2 :P

OpenStudy (alexwee123):

oh lol brain fart... :p x=y^2-4 x-4=y^2 y=(x-4)^1/2

OpenStudy (anonymous):

hehe it's cool. but (x+4)^(1/2) btw

OpenStudy (alexwee123):

lol :D i can't maf

OpenStudy (anonymous):

lol